OpenStudy (pphalke):

how many liters of a 50% acid solution must be mixed with a 15% acid solution to get 315 L of a 40% acid solution?

OpenStudy (aveline):

First, define your variables: A = liters of 50% acid solution B = liters of 15% acid solution A+B=320 liters Now what do we do?

OpenStudy (jdoe0001):

hmm

OpenStudy (jdoe0001):

any ideas on how to represent 50% and 15% and 40% as a decimal?

OpenStudy (jdoe0001):

and so

OpenStudy (pphalke):

@Aveline its 315 L you do .50+.015

OpenStudy (aveline):

Oops, 315 is correct. Make sure that you have the correct decimal place for 15% Here's the main equation: 0.50A+0.15B=(315)(0.40) Can you find a way to substitute B for A using this equation? A+B=315

OpenStudy (pphalke):

@Directrix

OpenStudy (aveline):

@pphalke Do you need me to walk you through it?

OpenStudy (aveline):

I'll write a detailed response. I can't reply after this because I'll be heading to dinner. A = liters of 50% acid solution B = liters of 15% acid solution These are the variables A+B=315 When you add the liters of A and B, they need to add up to 315 liters B=315-A Here, we are rearranging the previous equation. This way, we only have one unknown variable. 0.5A+0.15B=(315L)(0.40) This is the equation you will start off with. In words: one liter of a 50% acid solution plus one liter of a 15% acid solution equals 315 liters of a 50% acid solution 0.5A+0.15(315-A)=126 Substitute (315-A) for B. 0.5A+47.25-0.15A=126 Use the distributive property ...and simplify: 0.35A+47.25=126 0.35A=78.75 A=225 liters Remember the equation A+B=315? plug in A 225+B=315 B=90 liters A is 225 liters and B is 90 liters